# Can LLMs perform structured graph reasoning tasks?

### v3

@@ 2.1 Defining graph levels and complexity
  • Problem 1.3: Tree traversal with multiple possible solutions given a unweighted and undirected tree-based graph $G(n)$ of order $n$, the task involves finding the shortest path from node $v_{1}$ and $v_{n}$, among many possible paths from $v_{1}$ and $v_{n}$. There exists only one shortest possible path in this tree.
@@
  • Euler walk: Given a valid Euler graph $G$, an Euler path is defined as a path which covers all edges $E$ of the graph exactly once. The Eulerian path traversal is a more viable alternative to the Hamiltonian path traversal, which involves travelling all nodes $U$ of the graph exactly once, since the former can be evaluated computationally by a finite set of conditions, unlike the latter, which is an NP-complete problem and cannot be evaluated using a polynomial time algorithm. Based on this, we formulate the problem as follows.
@@
  • Checkpoint traversal This problem involves traversal of a tree-based graph $G$ between two nodes $u$ and $v$, such that a specific node $w$ must be part of the traversal path. Mathematically, this is equivalent to breaking down the traversal into two distinct problems – traversal from $u$ and $w$, and traversal from node $w$ and $v$. The goal of this cascaded traversal problem is to evaluate whether a single LLM response can solve more than one problem simultaneously. We hence formulate the following problem.
@@ 2.2 Graph generation, prompting and evaluation
Graphs, as well as corresponding solutions for all problems are automatically generated, in the form of adjacency matrices with nodes labelled alphabetically. We deliberately avoid adjacency lists (a more standard practice of computational representation), since by listing down the node connections, we make the problem somewhat easier for the LLM, thereby hindering our ability to assess the raw reasoning capabilities of the concerned models. We ensure that for all problems, there exists only a unique solution for all problems, so that evaluation can be unambiguous. For problems 3.1 (Euler walk), since there can be no guarantee of a single eulerian path, we convert the problem into a True/False problem. We also ensure that the distribution of valid and invalid euler graphs is balanced. All remaining problems require LLMs to provide a sequence of nodes in response. Thus, while lower LLM accuracies are expected, a more comprehensive evaluation can be performed. Across all problems, we generate three variations (or orders) of problems, depicting increasing levels of complexities.
@@
    O(20) jumbled: Unlike the previous two cases, where the labels of the nodes are more or less in order when traversed depth first from the starting node, we now jumble the labels of the nodes in O(20), to make the problem more complex. Hence, while the rows and columns of the adjacency matrix are still alphabetically ordered, the solution node sequence is completely randomized. This helps us evaluate whether LLMs rely on training biases such as order of nodes to return solutions.
@@ 3 Results and Analysis
Using the prompting and evaluation technique defined above, we evaluate the various models over the various problems. We evaluate the results of graph reasoning for problem 1, problem 2, and problem 3, by calculating averaged and normalized binary accuracies across 10 examples for each setting, and go from 0 to 1 in increments of 0.1. A model response is only marked correct if it predicted the shortest/least-cost traversal path correctly and completely. Along with binary accuracies, partial accuracies are also calculated. While we use binary accuracy to filter out models with weak reasoning abilities, we use partial accuracy for more granular comparison. These results can be visualized in Fig. 2. A detailed tabulation of these results are included in the Appendix, in tables 1 through 3.
@@
Figure 4: Comparison of graph order performance in various LLMs. LLM performance consistently drops as the order of the graph is increased. Also, while keeping the order of graph magnitude at 20, a drop is observed in performance when the nodes are jumbled, depicting a bias in models to expect node order to be alphabetically arranged.
@@ 3.2 Overall observations from Problem 1 and Problem 2
Although all models perform decently on trivial graph traversal tasks (e.g. problem 1.1), differentiation among models is quickly observed as more constraints are added. Referring to Fig 2, a general aggregate trend is observed where some reasoning abilities are observed in the case of GPT4, Claude-2, and GPT-3.5 (in this order of decreasing relative reasoning abilities). On the other hand, LLAMA-2 and Palm2 perform under 30% on the majority ($\geq$ 50% of the settings) of the tasks, depicting poor reasoning abilities. In more complex settings (such as O(20) graphs and O(20) jumbled graphs, not presented in this paper for simplicity and readability, but present in the code repository), an almost random-like performance is seen in these models. An anomaly is observed in problem 3.1 where a spike is observed in the performance of the aforementioned 2 models. However, since this particular problem is a True/False problem with a random baseline accuracy of 50%, the spike is rendered meaningless. Keeping this in mind, in the remainder of the paper, we primarily focus on GPT4, Claude-2, and GPT-3.5, since these contenders depict fairly reasonable graph reasoning abilities. However, all the results are openly available for analysis in our code repository.
@@ 3.3 Weighted vs unweighted graph traversal
Adding weights to edges of a graph in a least-cost traversal problem imposes a significant constraint, since it requires an additional memory component to track the sum of weights. Unsurprisingly, we observe a general trend across all models of dropping performance as soon as weights are added to trees (problems 1.3 vs 1.4) and well as grids (problems 2.1 vs 2.2). GPT-4 and Claude-2 demonstrate some ability to solve weighted traversal, demonstrating their innate ability to track numerical variable states in parallel to graph paths. On the other hand, chain-of-thought prompting [22] in GPT-3.5 demonstrates that the models tend to neglect weights and solve problems using an unweighted approach. The remainder of the models such as LLAMA-2 and PaLM 2 tend to respond with an algorithmic response rather than the solution to the specific problem posed. This is especially observed in 0-shot settings.
@@ 3.4 Effect of k-shot prompting
In Fig 3, we compare the effect of k-shot prompting across 8 tasks, and observe that in the majority of the tasks, k-shot prompting has either no statistically significant effect, or a negative effect on reasoning accuracy. In general, while few-shot prompting is helpful in response format shaping, it is of no particular aid for analytical tasks such as graph reasoning. Carefully examining individual responses, we observe models (even the more powerful ones like GPT4) confusion between different examples, and responding to k-shot examples rather than the relevant graph traversal question.
@@ 3.5 Effect of increased graph order and jumbling node order
In Fig 4, we compare the effect of different graph orders and complexities. In particular, we compare the O(10), O(20), and O(20) jumbled graph settings. We consistently observe across all models that a performance drop is observed when graph order is increased. An interesting insight enlightened when manually analyzing individual solutions is that poor performance in O(20) graphs in comparison to O(10) graphs is not only because of longer calculations, but because models tend to omit higher node labels (e.g. node labels J,K,L,M and so on, if all the nodes were alphabetically arranged), showing that the models were mostly trained on low-order graphs.
This behavior is amplified in the case of jumbled node sequences. Even linear graph traversal (an extremely trivial traversal problem) proves to be challenging to all models. This highlights a strong bias within models towards the expectation of nodes to appear in some form of order, even in few-shot settings. The only exception (in some cases) is GPT-4, which attempts to solve graphs through valid logical principles. However, this approach is again, majorly unsuccessful across grid-based traversals.
@@ 3.6 Positive response bias within LLMs
Through a special problem (problem 2.4), we observe that all models also consistently fail to recognize when no solution is possible. An interesting observation is that models fail to recognize when no valid solution is possible even in few-shot prompt settings, considering it is expected that this method should induce a formatting bias in the response. This indicates a clear training regime flaw among models to prefer avoiding empty responses at the cost of predicting wrong solutions. However, given that GPT-4 and Claude-2 do possess some valid logical reasoning and state tracking abilities, we observe some ability to predict the inexistence of any valid solution in these models. Surprisingly, in the case of few-shot prompt (where we prompt models with examples of valid graphs with solutions, the aforementioned models tend to identify the absence of a valid path with better accuracy (see Fig 3).
@@ 4 PathCompare prompting
Keeping these requirements in mind, we propose a novel prompting technique specifically for graph reasoning (traversal) tasks, which we refer to as the PathCompare technique. By simply adding the line “Let’s list down all the possible paths from node {X} to node {Y}, and compare the cost to get the answer.", we observe a significant improvement in the graph reasoning performance of all LLMs (see Figure 5 for visualization and Appendix Table 5 for detailed analysis), in comparison to standard prompting and other forms of prompting such as CoT (see Appendix Table 4). Through this prompt, we shape an LLM response to first list down multiple paths between two specified nodes. Once multiple paths are enumerated, the task boils down to a retrieval-cum-comparison task, which increases the probability of a more accurate response.
Figure 5: Comparison of different prompting techniques. Our proposed prompting technique (PathCompare) demonstrates an improvement in the majority of tasks across all models, i.e. 5/8 tasks for GPT3.5, 6/8 tasks for GPT4 and 5/8 tasks for Claude-2. However, one limitation of PathCompare is that it enhances positive response bias, as observed in the case of problem 2.4
@@ 5 Summary
Through this paper, we probe deeper into the graph-reasoning capabilities of various LLMs. We build on top of the work of [17], who highlight various trends related to graph reasoning, such as basic graph-reasoning abilities in LLMs, the lack of any improvement in performance of k-shot prompts, and various response biases. This study adopts a more granular approach towards analysing LLMs. First, we select a wider range of models from different model families (i.e. different architectures, training data and methods and distinct fine-tuning methods), and systematically compare the different models over reasoning tasks. Our problem design evaluates these models over increasingly complex problems and added constraints. Through this process, we highlight some interesting properties of LLMs. First off, we highlight various type of constraints that lead to performance drop (such as weighted edges and unordered node sequences). We also conclude that k-shot prompts are unhelpful for analytical tasks such as graph traversal. We also highlight how various models are biased towards providing a positive response. This, hence, leads to these LLMs providing wrong responses, even when no valid solution exists. Secondly, we highlight the root of poor graph-based reasoning in LLMs, and link this to the inability to backtrack solutions to compare and come to an optimal conclusion. Building on top of this, our novel prompting technique PathCompare helps mitigate this issue and improve graph reasoning abilities in LLMs.

### v4

@@ 2.1 Defining graph levels and complexity
  • Problem 1.3: Tree traversal with multiple possible solutions: given a unweighted and undirected tree-based graph $G(n)$ of order $n$, the task involves finding the shortest path from node $v_{1}$ to $v_{n}$, among many possible paths. There exists only one shortest possible path in this tree.
@@
  • Euler walk: Given a valid Euler graph $G$, an Euler path is defined as a path which covers all edges $E$ of the graph exactly once. The Eulerian path traversal is a more viable alternative to the Hamiltonian path traversal, which involves travelling all nodes $U$ of the graph exactly once, since the former can be evaluated computationally by a finite set of conditions, unlike the latter, which is an NP-complete problem and cannot be evaluated using a polynomial time algorithm. Hence, we formulate the problem as follows.
@@
  • Checkpoint traversal This problem involves traversal of a tree-based graph $G$ between two nodes $u$ and $v$, such that a specific node $w$ must be part of the traversal path. Mathematically, this is equivalent to breaking down the traversal into two distinct problems – traversal from $u$ and $w$, and traversal from node $w$ and $v$. The goal of this cascaded traversal problem is to evaluate whether a single LLM response can simultaneously solve multiple problems. We hence formulate the following problem.
@@ 2.2 Graph generation, prompting and evaluation
Graphs and their corresponding solutions are automatically generated as adjacency matrices with alphabetically labeled nodes. We deliberately avoid adjacency lists (a more standard practice of computational representation), to avoid triviality, since a direct listing of individual node connections simplifies the complexity for the LLM, thereby hindering the ability to assess its raw reasoning capabilities. We ensure that there exists only one unique solution for all problems, so that evaluation can be unambiguous. For problem 3.1 (Euler walk), we convert it into a True/False problem due to the lack of a guaranteed single Eulerian path, ensuring a balanced distribution of valid and invalid cases. All remaining problems require LLMs to provide a sequence of nodes in response. Thus, while lower LLM accuracies are expected, a more comprehensive evaluation can be performed. Across all problems, we generate three variations (or orders) of problems, depicting increasing levels of complexities.
@@
    O(20) jumbled: In contrast to the previous cases, where node labels follow a near-ordered sequence in depth-first traversal, the labels of the nodes in O(20) are randomized to increase complexity. Although the adjacency matrix remains alphabetically ordered, the solution sequence is entirely shuffled. This approach allows us to assess whether LLMs depend on training biases, such as node order, when generating solutions.
@@ 3 Results and Analysis
Using the prompting and evaluation technique defined above, we evaluate models over the defined problems. We evaluate the results of graph reasoning for all problems by calculating averaged and normalized binary accuracies across 10 examples for each setting, and go from 0 to 1 in increments of 0.1. A model response is only marked correct if it predicted the shortest/least-cost traversal path correctly and completely. Along with binary accuracies, partial accuracies are also calculated. While we use binary accuracy to filter out models with weak reasoning abilities, we use partial accuracy for more granular comparison. These results can be visualized in Fig. 2. A detailed tabulation of these results are included in the Appendix tables 1 through 3.
@@
Figure 4: Comparison of graph order performance in various LLMs. LLM performance accuracy consistently drops as the order of the graph is increased. Also, while keeping the order of graph magnitude at 20, a drop is observed in performance when the nodes are jumbled, depicting a bias in models to expect node order to be alphabetically arranged.
@@ 3.2 Overall observations from Problem 1 and Problem 2
Although all models perform decently on trivial graph traversal tasks (e.g. problem 1.1), differentiation among models is quickly observed as more constraints are added. Referring to Fig. 2, a general aggregate trend is observed where some reasoning abilities are observed in the case of GPT4, Claude-2, and GPT-3.5 (in this order of decreasing relative reasoning abilities). On the other hand, Llama-2 and Palm2 perform under 30% on the majority ($\geq$ 50% of the settings) of the tasks, depicting poor reasoning abilities. In more complex settings (such as O(20) graphs and O(20) jumbled graphs, not presented in this paper for simplicity and readability, but present in the code repository), an almost random-like performance is seen in these models. An anomaly is observed in problem 3.1 where a spike is observed in the performance of the aforementioned 2 models. However, since this particular problem is a True/False problem with a random baseline accuracy of 50%, the spike is rendered meaningless. Keeping this in mind, in the remainder of the paper, we primarily focus on GPT4, Claude-2, and GPT-3.5, since these contenders depict fairly reasonable graph reasoning abilities. However, all the results are openly available for analysis in our code repository.
@@ 3.3 Weighted vs unweighted graph traversal
Adding weights to graph edges in least-cost traversal problems introduces significant complexity, requiring models to track cumulative weights. Unsurprisingly, we observe a general trend of dropping performance as soon as weights are added to trees (problems 1.3 vs 1.4) and well as grids (problems 2.1 vs 2.2). GPT-4 and Claude-2 demonstrate some ability to solve weighted traversal, demonstrating their innate ability to track numerical variable states in parallel to graph paths. On the other hand, chain-of-thought prompting [22] in GPT-3.5 demonstrates that the models tend to neglect weights and solve problems using an unweighted approach. The remainder of the models such as Llama-2 and PaLM 2 tend to respond with a description of the algorithm underlying the solution, rather than the solution itself. This is especially observed in 0-shot settings.
@@ 3.4 Effect of k-shot prompting
In Fig. 3, we compare the effect of k-shot prompting across 8 tasks, and observe that in the majority of the tasks, k-shot prompting has either no statistically significant effect, or a negative effect on reasoning accuracy. In general, while few-shot prompting is helpful in response format shaping, it is of no particular aid for analytical tasks such as graph reasoning. Carefully examining individual responses, we observe models (even the more powerful ones like GPT4) confusion between different examples, and responding to k-shot examples rather than the relevant graph traversal question.
@@ 3.5 Effect of increased graph order and jumbling node order
In Fig. 4,We analyze the impact of different graph orders and complexities, specifically comparing O(10), O(20), and O(20) jumbled graph settings. We consistently observe across all models that a performance drop is observed when graph order is increased. An interesting insight enlightened when manually analyzing individual solutions is that poor performance in O(20) graphs in comparison to O(10) graphs is not solely due to increased computational demands but also because models often omit higher node labels (e.g., J, K, L, M and so on) when labels are alphabetically ordered, indicating training on lower-order graphs.
This behavior is amplified in the case of jumbled node sequences, where even trivial linear traversal becomes challenging, highlighting a strong model bias towards ordered node sequences, even in few-shot settings. GPT-4, while occasionally applying valid logical principles, still struggles across grid-based traversals.
@@ 3.6 Positive response bias within LLMs
Through a special problem (problem 2.4), we observe that all models also consistently fail to recognize when no solution is possible. An interesting observation is that models fail to recognize when no valid solution is possible even in few-shot prompt settings, considering it is expected that this method should induce a formatting bias in the response. This indicates a clear training regime flaw among models to prefer avoiding empty responses at the cost of predicting wrong solutions. However, given that GPT-4 and Claude-2 do possess some valid logical reasoning and state tracking abilities, we observe some ability to predict the inexistence of any valid solution in these models. Surprisingly, in the case of few-shot prompt (where we prompt models with examples of valid graphs with solutions, the aforementioned models tend to identify the absence of a valid path with better accuracy (see Fig. 3).
@@ 4 PathCompare prompting
Keeping these requirements in mind, we propose a novel prompting technique specifically for graph reasoning (traversal) tasks, which we refer to as the PathCompare technique. By simply adding the line “Let’s list down all the possible paths from node {X} to node {Y}, and compare the cost to get the answer.", we observe a significant improvement in the graph reasoning performance of all LLMs (see Fig. 5 for visualization and Appendix Table 5 for detailed analysis), in comparison to standard prompting and other forms of prompting such as CoT (see Appendix Table 4). Through this prompt, we shape an LLM response to first list down multiple paths between two specified nodes. Once multiple paths are enumerated, the task boils down to a retrieval-cum-comparison task, which increases the probability of a more accurate response.
Figure 5: Comparison of different prompting techniques. Our proposed prompting technique (PathCompare) demonstrates an improvement of accuracy in the majority of tasks across all models, i.e. 5/8 tasks for GPT3.5, 6/8 tasks for GPT4 and 5/8 tasks for Claude-2. However, one limitation of PathCompare is that it enhances positive response bias, as observed in the case of problem 2.4.
@@ 5 Summary
Through this paper, we probe deeper into the graph-reasoning capabilities of various LLMs. This study adopts a more granular approach towards analysing LLMs, in comparison to previous works [17]. First, we select a wider range of models from different model families (i.e. different architectures, training data and methods and distinct fine-tuning methods), and systematically compare the different models over reasoning tasks. Our problem design evaluates these models over increasingly complex problems and added constraints. Through this process, we highlight some interesting properties of LLMs. First off, we highlight various type of constraints that lead to performance drop (such as weighted edges and unordered node sequences). We also conclude that k-shot prompts are unhelpful for analytical tasks such as graph traversal. We also highlight how various models are biased towards providing a positive response. This, hence, leads to these LLMs providing wrong responses, even when no valid solution exists. Secondly, we highlight the root of poor graph-based reasoning in LLMs, and link this to the inability to backtrack solutions to compare and come to an optimal conclusion. Building on top of this, our novel prompting technique PathCompare helps mitigate this issue and improve graph reasoning abilities in LLMs.
@@ A  B  C  D  E  F  G  H  I  J
A more visually effective representation of the matrix is presented below, for the understanding of the users, although the above plain text version is used for prompting the LLMs. The numbers in red depict the connections that constitute the solution to the corresponding graph traversal.
 A  B  C  D  E  F  G  H  I  J
 A  0  1  0  0  0  0  0  0  0  0
 B  0  0  1  0  0  0  0  0  0  0
 C  0  0  0  1  0  0  0  0  0  0
 D  0  0  0  0  1  0  0  0  0  0
 E  0  0  0  0  0  1  0  0  0  0
 F  0  0  0  0  0  0  1  0  0  0
 G  0  0  0  0  0  0  0  1  0  0
 H  0  0  0  0  0  0  0  0  1  0
 I  0  0  0  0  0  0  0  0  0  1
 J  0  0  0  0  0  0  0  0  0  0